OpenStudy (australopithecus):

let z^3 = -8i Find all z in the form a + bi I need someone to tell me if I'm doing this correctly, I got r = 8, theta = 3pi/2 = arcsin(-8/8) = arcsin(-1) Then I got the roots, 1 + (3)^(1/2)i 2 1-(3)^(1/2)i Did I make a mistake? Should I have set theta to -pi/2 instead of 3pi/2 I'm kind of confused in regards to that

OpenStudy (nipunmalhotra93):

r=2.

OpenStudy (australopithecus):

r = (-8^2)^(1/2) = (64)^(1/2) = 8

OpenStudy (australopithecus):

I dont think you are right in regards to r = 2

OpenStudy (nipunmalhotra93):

it's r^3=8

OpenStudy (australopithecus):

oh yeah I know

OpenStudy (australopithecus):

I gave the roots I got

OpenStudy (australopithecus):

my general solution is, \[z_0 = 2(\cos(\frac{\pi}{3} + \frac{2\pi k}{3}) + \sin(\frac{\pi}{3} + \frac{2\pi k}{3})i)\] where, k = 0, 1, 2

OpenStudy (anonymous):

Still need help?

OpenStudy (australopithecus):

My question is should I have used -pi/2 instead of 3pi/2 they both equal the same thing but does it make a difference in regards to this question are my roots correct?

OpenStudy (anonymous):

One moment, I'm going to check

OpenStudy (australopithecus):

Is there a way to check if your roots are correct make sense?

OpenStudy (nipunmalhotra93):

@Australopithecus shouldn't it be Pi/2 instead of Pi/3 in your answer?

OpenStudy (australopithecus):

thank you I really appreciate the help

OpenStudy (anonymous):

I just finished the problem. I'm going to attach the file.

OpenStudy (anonymous):

http://i.imgur.com/Dkf54kX.jpg

OpenStudy (anonymous):

You're welcome! I hope you understand it :) Good luck!

OpenStudy (nipunmalhotra93):

@Australopithecus The angle you choose has to be in between 0 and 2Pi. So 3Pi/2 is the right choice.

OpenStudy (australopithecus):

well, (3pi/2)*1/3 = pi/2

OpenStudy (australopithecus):

good eye lol :) yeah you are rght nipunmalhotra93

OpenStudy (australopithecus):

thanks nipunmalho and jesstho wish I could give multiple medals :S

OpenStudy (anonymous):

Lol you're welcome